OpenStudy (anonymous):

swimming pool is 1500 meters. length is 10meters less than width what is width

OpenStudy (anonymous):

I'm assuming the area of the swimming pool is 1500 meters squared?

OpenStudy (anonymous):

yes 1500 meters square

OpenStudy (anonymous):

L=W-10. W*L=1500. W(W-10)=1500. W^2-10W-1500=0. See if you can do it from here.